i want to solve thisimultaneous equation 5x-5y=5,4x-3y=25
i took the first equation x=5+5y/5 and them
4(5-5y)/5-3y=25
i donīt understand the siplpification

Found 2 solutions by checkley77, ankor@dixie-net.com:
Answer by checkley77(12844)

5x-5y=5 multiply this equation by -4
4x-3y=25 multiply this equation by 5 & then add them
-20x+20y=-20
20x-15y=125
-----------
5y=105
y=105/5
y=21 ans.
5x-5*21=5
5x-105=5
5x=5+105
5x=110
x=110/5
x=22 ans.
Proof:
4*22-3*21=25
88-63=25
25=25


Answer by ankor@dixie-net.com(22740)

solve these simultaneous equations:
5x - 5y = 5
4x - 3y = 25
:
Take the 1st equation:
5x - 5y = 5
Simplify, divide by 5:
x - y = 1
x = y + 1; use this for substitution for x in the 2nd equation:
:
4x - 3y = 25
4(y+1) - 3y = 25
4y + 4 - 3y = 25
4y - 3y = 25 - 4
y = 21
:
Find x:
x = y + 1
x = 21 + 1
x = 22
:
:
Check solution in the 2nd equation
4(22) - 3(21) = 25
88 - 63 = 25; confirms our solutions